OpenStudy (simplymarie_x):

Factor these equations: x^2 – 7x + 10 4x^2 – 81 p (x) = 3x^3 – 12x

rishavraj (rishavraj):

for x^2 -7x + 10 = 0 clue : -5 - 2 = -7 (-5)*(-2) = 10